OpenStudy (anonymous):

I WILL FAN AND MEDAL THE PERSON WHO HELPS ANSWER THE QUESTION Explain, in your own words, the step-by-step process to simplifying the expression below. Include the simplified answer in your explanation. (6x2 − 5) − (3x2 + x + 7)

OpenStudy (anonymous):

First we must work in the brackets. As per convention we'll start with the bracket on the left. (6x2 - 5) The first thing to do is deal with the multiplication 6x2 = 12 Next we subtract (12-5) =7 Thus the left bracket is 7, now we move to the right bracket. Same thing we deal with the multiplication which is 3x2 (6+x+7) Using the properties of addition we can add 6 and 7 which gives us 13 So now the line looks like this (7) - (13+x) Distribute the negative over the 13 and the x and it'll be 7 - 13 - x To isolate for x set the entire line to 0 7 - 13 - x = 0 -6 - x = 0 x = 6. Hopefully that answers your question, and it wasn't 6x^2 and 3x^2 in the brackets. Haha then the process is really different.

OpenStudy (anonymous):

Excellent thanks so much @SA5UK3